a man mixes 100 lbs of 90% copper alloy with 150 lbs of 60% copper alloy what is his percentage of the total copper alloy?
Answer by stanbon(75887)

a man mixes 100 lbs of 90% copper alloy with 150 lbs of 60% copper alloy what is his percentage of the total copper alloy?
------------------
Equation:
copper + copper = copper
0.90*100 + 0.60*150 = x(100+150)
Multiply thru by 100 to get:
90*100 + 60*150 = x(25000)
25000x = 9000+9000
x = 0.72 = 72%
====================
Cheers,
Stan H.
====================